# subequation

### v1

@@ 1 Introduction
In the past two decades we have seen the discovery of numerous heavy exotic hadrons Brambilla:2019esw . These are states that do not fall in the traditional quark model classification of mesons and baryons as $q\bar{q}$ and $qqq$ states, respectively. In particular, the majority of the exotic heavy hadrons have been discovered in the charmonium and bottomonium spectra close and above the heavy meson-antimeson thresholds. The lack of a comprehensive and QCD based description of these states shows a significant gap in our understanding of the strong interactions.
@@
Beyond the quenched approximation the spectrum of static energies becomes populated by additional states such as heavy meson-antimeson pairs and states containing light-quark hadrons. These open the possibility of transitions and decays for conventional and hybrid quarkonium. Experimental measurements show that some of the states above open flavor thresholds have significant branching ratios for decays to heavy meson-antimeson pairs, however there are some intriguing exceptions. Moreover, some of the exotic states appear to be extremely close to the heavy meson thresholds, which has prompted their interpretation as heavy-meson shallow bound states Braaten:2003he ; Esposito:2016noz ; Guo:2017jvc . Therefore, it is crucial to elucidate the couplings of the Born-Oppenheimer states to the open flavor thresholds and to determine if they are narrow enough to be observed.
@@
A very related problem is that of the conventional quarkonium mixing with and decaying to heavy meson-antimeson pairs. This problem has been studied in Refs. Bicudo:2019ymo ; Bicudo:2020qhp ; Bicudo:2022ihz ; Bruschini:2020voj ; Bruschini:2021sjh ; Bruschini:2023zkb using models based on an extension of the Born-Oppenheimer approximation, as well as using BOEFT in Ref. TarrusCastella:2022rxb . The present article can be considered as a continuation of the latter.
@@ 2 Continuum states in BOEFT
Similarly, we represent the heavy hadron pairs by ${\cal M}_{\kappa\alpha}(t,\,\bm{r},\,\bm{R})$. In this case $\kappa$ is any of the possible combinations of the light degrees of freedom spin of each one of the hadrons forming the pair. Note, that each one of these pairs may be mapped to more than one field ${\cal M}_{\kappa}$, which in turn, for a specific heavy-quark spin state, will correspond to a combination of the physical heavy hadron pair states. Both the $\Psi$ and ${\cal M}$ fields carry light-quark flavor quantum numbers, which for the purpose of this work must be equal. We do not show them explicitly since they do not affect the construction of the EFT. However, it should be kept in mind that the matching coefficients, such as the potentials, do depend on them.
@@
$\Psi_{\kappa\alpha}$ $=\sum_{\lambda}\left(P_{\kappa\lambda}(\hat{\bm{r}})\right)_{\alpha}\Psi_{\kappa\lambda}\,,$ (3)
@@
In the $D_{\infty h}$ basis in Eqs. (3)-(4) the static potential $V_{\Psi_{\kappa}}^{(0)}$ is diagonal Brambilla:2017uyf , however, due to the dependence of the projection vectors on the angles of $\hat{\bm{r}}$, in this basis the kinetic operator becomes non-diagonal. These off-diagonal terms are the so-called nonadiabatic coupling which induces a mixing between the $\Psi_{\kappa\lambda}$ components. A more in-depth discussion can be found in Refs. Berwein:2015vca ; Brambilla:2017uyf ; Oncala:2017hop ; Soto:2020pfa .
The static potential between the heavy hadron pairs is not known in general. The static energies between heavy meson-meson and heavy meson-antimeson pairs has been studied on the lattice in Refs. Bali:2005fu ; Bicudo:2012qt ; Brown:2012tm ; Bicudo:2015vta ; Bicudo:2015kna ; Prelovsek:2019ywc ; Bulava:2019iut ; Bicudo:2021qxj ; Sadl:2021bme . The results show that the static energies are mostly flat lines at the energy corresponding to the heavy meson masses except in some cases in the short-distance limit where the potential is attractive or repulsive depending on the specific heavy meson pair. However, only in Refs. Bali:2005fu ; Bulava:2019iut the coupled system of the heavy meson-antimeson pair and quarkonium static energies has been studied. The analysis of Ref. TarrusCastella:2022rxb indicates that when the coupling is taken into account the static potential of the heavy meson-antimeson pair becomes flat. Hence, we are going to assume that
@@
with $\Psi^{(n\ell)}_{\kappa\alpha}$ the wave function solution of the following coupled Schrödinger equations
@@
with $n$ the principal quantum number and the components of the tensor spherical harmonics Khersonskii:1988krb defined as
@@
For conventional quarkonium coupled to the lowest lying heavy meson-antimeson pair we recover [^3] To match the results of Ref. TarrusCastella:2022rxb a $(-1)^{\kappa^{\prime}}$ phase has to be shifted from the definition of Eq. (19) to Eq. (20). the results of Ref. TarrusCastella:2022rxb .
Now we can compute the continuum states contribution to the double-heavy hadron self-energy
@@
where $p(E)=\sqrt{2\mu(E+i\epsilon+2m_{Q}-m_{T})}$ with $\mu$ and $m_{T}$ being the heavy hadron pair reduced and total masses, respectively. In order to take into account the self-energy to all orders in perturbation theory, we can resume the self-energy into the double-heavy hadron propagator. Using Sokhotsky’s formula we separate the real and imaginary parts of the self-energy and find the bound state energy of the double-heavy hadron as the roots of
@@
For the rest of the paper we will use $E_{n\ell}=E^{(0)}_{n\ell}$. However, one of the strengths of our approach is that we can improve the accuracy of this quantity by including higher order terms in BOEFT. This is important, because when $E_{n\ell}+2m_{Q}-m_{T}$ is close to zero, i.e when the double-heavy hadron state is close to the heavy hadron pair threshold, the contribution of the continuum states is very sensitive to small variations of $E_{n\ell}$.
3 Hybrid quarkonium and heavy meson pairs
Let us now apply the results of the previous section to study the effect of the heavy meson-antimeson thresholds to the hybrid quarkonium spectrum. In table 1 we show the quantum numbers of four lowest mass gluelumps Herr:2023xwg and the $D_{\infty h}$ representations of the corresponding static states Brambilla:1999xf ; Berwein:2015vca . Each heavy meson-antimeson pair is characterized by the spin and parity of the light-quark states forming the two heavy mesons, which we label as $\kappa_{1}^{p_{1}}$ and $\kappa_{2}^{p_{2}}$ in Table 2. Combining the spin and parity of the light quarks we find their total $\kappa^{pc}$. From these, we work out the possible $\Lambda^{\sigma}_{\eta}$ representations in Table 2. A hybrid (or conventional) quarkonium state couples to a heavy meson-antimeson pair if they can be projected into the same $D_{\infty h}$ representation. For the lowest lying hybrid quarkonium, with $\kappa=1^{+-}$ gluelump, the only matching $D_{\infty h}$ representation with the lowest lying heavy meson-antimeson pairs, i.e. $D_{(s)}^{(*)}\bar{D}_{(s)}^{(*)}(B_{(s)}^{(*)}\bar{B}_{(s)}^{(*)})$, is $\Sigma_{u}^{-}$ corresponding to a $0^{-+}$ light-quark state. For this case the form factor in Eq. (20) takes the following simple form
@@ 3 Hybrid quarkonium and heavy meson pairs
Table 2: Total spin ($\kappa$), parity ($p$) and charge conjugation ($c$) of the light-quarks states and pair combinations for the three lightest light-quark states forming heavy mesons and the corresponding possible projections into $D_{\infty h}$ representations. Each block of states, separated by a single horizontal line, corresponds to degenerate or nearly degenerate states. Note that the representation of the light-antiquark state should be conjugated.
We can infer the following selection rules for the decays of hybrid quarkonium into $D_{(s)}^{(*)}\bar{D}_{(s)}^{(*)}(B_{(s)}^{(*)}\bar{B}_{(s)}^{(*)})$ from Eq. (25). First, we can see that only states with a $\lambda=0$ component can decay through this channel [^4] This has been recently noted in Ref. Bruschini:2023tmm .. In table 3 we summarize the details of the $\kappa=1^{+-}$ gluelump hybrid quarkonium states Berwein:2015vca ; Oncala:2017hop . Since the $\lambda=0$ component corresponds to the $\Sigma_{u}^{-}$ static state, we can read off table 3 which hybrid states are allowed to decay through this channel. Using the notation $l_{\ell}$, these are $s_{0}$, $(s\backslash d)_{1}$ and $(p\backslash f)_{2}$. On the other hand the states $p_{1}$ and $d_{2}$ are forbidden to decay. Higher angular momentum, not shown in table 3, with mixed $l$ are also allowed to decay. It is interesting to note that the mixed states are dominated by the $\Pi_{u}$ component Berwein:2015vca , which adds a small suppression to the decays into heavy mesons. The second selection rule is that the partial wave of the heavy meson pair must be equal to the total angular momentum eigenvalue of the hybrid quarkonium states $\ell$. This is a consequence of the light-quark state of the heavy-meson pairs coupling to hybrids being spin 00. Therefore, only one partial wave is allowed in the decays of quarkonium hybrids into $D^{(*)}\bar{D}^{(*)}(B^{(*)}\bar{B}^{(*)})$ at leading order in BOEFT.
In previous works Tanimoto:1982eh ; LeYaouanc:1984gh ; Iddir:1988jc ; Ishida:1991mx ; Close:1994hc ; Page:1996rj ; Page:1998gz ; Kou:2005gt , it had been argued that quarkonium hybrids could not decay into heavy-meson pairs in $S$-wave, our results show that in the Born-Oppenheimer approach this is not strictly true, since S-wave decays are possible, but at leading order only for $\ell=0$ quarkonium hybrids.
@@ 00 $s$ $\Sigma_{u}^{-}$$\{0^{++},1^{+-}\}$ $H_{3}$
Table 3: Quantum numbers of hybrid quarkonium with $1^{+-}$ gluelump. The quantum numbers are as follows: $l(l+1)$ is the eigenvalue of $\bm{L}^{2}_{Q\bar{Q}}$, the square of the angular momentum of the heavy-quark pair. $\ell(\ell+1)$ is the eigenvalue of $\bm{L}^{2}=(\bm{L}_{Q\bar{Q}}+\bm{S}_{1})^{2}$, with $\bm{S}_{1}$ the gluelump spin operator. $s_{Q\bar{Q}}(s_{Q\bar{Q}}+1)$ is the eigenvalue of $\bm{S}^{2}_{Q\bar{Q}}$, the square of the total heavy-quark spin operator. The square of the total spin $\bm{J}^{2}=(\bm{L}+\bm{S}_{Q\bar{Q}})^{2}$ has eigenvalue $j(j+1)$. As we are working at leading order in the heavy-quark mass expansion the hybrid quarkonium masses only depend on $\ell$ and a principal quantum number $n$. The set of total spin states for each $\ell$ form degenerate multiplets at this order. The $\Lambda^{\sigma}_{\eta}$ column indicates the $D_{\infty h}$ representations of static energies that contribute to states. For $\ell>0$, two sets of solutions with opposite parity exists. One of these corresponds to the mixing the $\Sigma_{u}^{-}$ and $\Pi_{u}$ static states while the other is purely $\Pi_{u}$erwein:2015vca . The solutions that mix static states also mix the heavy-quark pair angular momentum Oncala:2017hop . The last column indicates the labels for the spin-symmetry multiplets used in Refs. Braaten:2014qka ; Berwein:2015vca .
@@
Figure 1: Leading order diagrams in weakly-coupled pNQRCD contributing to the hybrid quarkonium and heavy meson pair coupling potential. The diagrams are in coordinate space. The single, double and curly lines represent the heavy-quark singlet, heavy-quark octet and gluon fields respectively. The dashed line corresponds to the light-quark fields. The black dots stand for the dipolar chromoelectric operators. The circles with a cross and with a diamond represent the quarkonium hybrid and heavy meson-antimeson pair states, respectively. It should be noted that while at leading order the quarkonium hybrid only overlaps with the octet field, the heavy meson pair can overlap both with the singlet and octet fields. The gray circles represent nonperturbative dynamics of the light degrees of freedom.
The short-distance region is, by construction, the region where the multipole expansion is valid. Therefore, in this region the potential can be expanded in a series in powers of $r$. Such series can be obtained directly using weakly-coupled pNRCQD Pineda:1997bj ; Brambilla:1999xf , which is an EFT incorporating the heavy-quark mass and multipole expansions. Several similar examples can be found in Refs. Oncala:2017hop ; Brambilla:2019jfi ; Soto:2021cgk . This computation yields the coefficients of the expanded potential as correlators of gluon and/or light-quark operators, which encode the nonperturbative dynamics. The two weakly-coupled pNRCQD leading order diagrams contributing to mixing potential are depicted in Fig. 1. We must consider two diagrams because the heavy meason-antimeson pair state can overlap both with the heavy-quark singlet and octet fields. However, in both instances, a single chromoelectric dipole coupling in enough to produce the transition between quarkonium hybrid and heavy meson pair. Therefore, at leading order the short-distance coupling potential must be proportional to $r$.
@@
The long distance quarkonium-heavy meson pair mixing potential obtained from lattice data was to found behave as $n_{l}=3$ TarrusCastella:2022rxb .
@@
$V_{01\Sigma}(r)=w_{s}(r)V^{\,{\rm(s.d.)}}_{01\Sigma}(r)+w_{l}(r)V^{\,{\rm(l.d.)}}_{01\Sigma}(r)\,.$ (29)
with the interpolating functions $w_{s}=(r_{0}/(r+r_{0}))^{n_{l}+2}$ and $w_{l}=(r/(r+r_{0}))^{n_{l}+2}$ and with $n_{l}$ chosen as the smallest one with the correct short and long-distance limits. The $r_{0}$ parameter determines the value of $r$ where both interpolating functions are equal. We will use values of $r_{0}\sim 0.25$ fm since it was found in Ref. TarrusCastella:2022rxb to provide a good description of the quarkonium-heavy meson pair mixing potential obtained from the lattice data of Ref. Bali:2005fu as the lattice mixing potential reaches its maximum absolute value for distances $r\sim 0.25$ fm. Moreover, it is a reasonable value for the distance in which the multipole expansion breaks down and thus for the transition from the short-distance regime to the long-distance one.
We consider three different variations of the model potential. Models (i) and (ii) take the parameter values given by dimensional analysis in Eq. (28) and $n_{l}=3$ and $n_{l}=2$, respectively. Finally, model (iii) is tunned to roughly match the lattice quarkonium-heavy meson pair mixing potential to serve as a reference point. Its parameters read as $c_{s}=8\Lambda^{2}_{\rm QCD}$, $c_{l}=0.9\Lambda^{-2}_{\rm QCD}$, $n_{l}=3$ and $r_{0}=0.275$ fm. In Fig. 2 we plot the three model potentials and the quarkonium-heavy-meson pair mixing potential fitted in Ref. TarrusCastella:2022rxb to the lattice data of Ref. Bali:2005fu . The heavy meson-antimeson pair interpolating operator in Ref. Bali:2005fu is isospin $I=0$ corresponding to the normalized sum of the charged and neutral heavy meson-antimeson pairs. As we use this data as reference, we should also take ${\cal M}_{0^{-+}}$ to be $I=0$ and therefore the results obtained from this model potential will correspond to the sum of neutral and charged heavy meson pair contributions. Since the light-quarks in Ref. Bali:2005fu are degenerate it is also possible to take the field ${\cal M}_{0^{-+}}$ as carrying a single $q\bar{q}$ light-quark flavor and rescale the potential by $1/\sqrt{2}$. We can make use of this to obtain the coupling potential to $D_{s}^{(*)}\bar{D}_{s}^{(*)}(B_{s}^{(*)}\bar{B}_{s}^{(*)})$ in the $SU(3)$ flavor symmetry limit.
Figure 2: Plot of the three models for the quarkonium hybrid and heavy meson pair coupling potential. The short-distance parametrization is $\sim r$ as determined from weakly-coupled pNRQCD. In the first two models the long-distance pieces behave as $r^{-3}$ and $r^{-2}$, respectively, while all the parameters values are determined by dimensional analysis. In the third model the parameters are tunned, within natural sizes, to match the quarkonium and heavy meson-antimeson pair coupling potential obtained from the lattice computation of Ref. Bali:2005fu , which we use as a reference for our models. The latter corresponds to the line plotted in light gray.

### v2

@@ 1 Introduction
In the past two decades we have seen the discovery of numerous heavy exotic hadrons Brambilla:2019esw . These states do not fall in the traditional quark model classification of mesons and baryons as $q\bar{q}$ and $qqq$ states, respectively. A large amount of exotic heavy hadrons have been discovered in the charmonium and bottomonium spectra close and above the heavy meson-antimeson thresholds. The lack of a comprehensive and QCD based description of these states shows a significant gap in our understanding of the strong interactions.
@@
Beyond the quenched approximation, the spectrum of static energies becomes populated by additional states such as heavy meson-antimeson pairs and states containing light-quark hadrons. These open the possibility of transitions and decays for conventional and hybrid quarkonium. Experimental measurements show that some of the states above open flavor thresholds have significant branching ratios for decays to heavy meson-antimeson pairs, however there are some intriguing exceptions. Moreover, some of the exotic states appear to be extremely close to heavy meson thresholds, which has prompted their interpretation as heavy-meson shallow bound states Braaten:2003he ; Esposito:2016noz ; Guo:2017jvc . Therefore, it is crucial to elucidate the couplings of the Born-Oppenheimer states to the open flavor thresholds and to determine if they are narrow enough to be observed.
@@
A closely related problem is that of conventional quarkonium mixing with and decaying to heavy meson-antimeson pairs. This problem has been studied in Refs. Bicudo:2019ymo ; Bicudo:2020qhp ; Bicudo:2022ihz ; Bruschini:2020voj ; Bruschini:2021sjh ; Bruschini:2023zkb using models based on an extension of the Born-Oppenheimer approximation, as well as using BOEFT in Ref. TarrusCastella:2022rxb . In the present article we generalize the formalism presented in Ref. TarrusCastella:2022rxb to the coupling of double-heavy hadrons to pairs of heavy hadrons in Section 2. As an example of our formalism, in Section 3, we study the specific case of quarkonium hybrids coupled to heavy meson-antimeson pairs.
@@ 2 Continuum states in BOEFT
Similarly, we represent the heavy hadron pairs by ${\cal M}_{\kappa\alpha}(t,\,\bm{r},\,\bm{R})$. In this case $\kappa$ is any of the possible combinations of the light degrees of freedom spin of each one of the hadrons forming the pair. Note that, when the heavy quark spin is taken into account, a ${\cal M}_{\kappa}$ field can be decomposed into a linear combination of physical heavy hadron pair states and that each one of these pairs can appear in more than one field. Both the $\Psi$ and ${\cal M}$ fields carry light-quark flavor quantum numbers, which for the purpose of this work must be equal. We do not show them explicitly since they do not affect the construction of the EFT. However, it should be kept in mind that the matching coefficients, such as the potentials, do depend on them.
@@
$\Psi_{\kappa\alpha}$ $=\sum_{\lambda=-\kappa,..,\kappa}\left(P_{\kappa\lambda}(\hat{\bm{r}})\right)_{\alpha}\Psi_{\kappa\lambda}\,,$ (3)
@@
In the $D_{\infty h}$ basis in Eqs. (3)-(4) the static potential $V_{\Psi_{\kappa}}^{(0)}$ is diagonal Brambilla:2017uyf , however, due to the dependence of the projection vectors on the angles of $\hat{\bm{r}}$, in this basis the kinetic operator becomes non-diagonal. These off-diagonal terms are the so-called nonadiabatic coupling which induces a mixing between the $\Psi_{\kappa\lambda}$ components. A more in-depth discussion can be found in Refs. Berwein:2015vca ; Brambilla:2017uyf ; Oncala:2017hop ; Soto:2020pfa . In Refs. Bruschini:2020voj ; Bruschini:2021sjh ; Bruschini:2023zkb ; Bruschini:2023tmm , the basis in terms of spherical spin components is referred as diabatic basis, while the $D_{\infty h}$ basis is called adiabatic basis.
The static potential between the heavy hadron pairs is not known in general. The static energies between heavy meson-meson and heavy meson-antimeson pairs has been studied on the lattice in Refs. Bali:2005fu ; Bicudo:2012qt ; Brown:2012tm ; Bicudo:2015vta ; Bicudo:2015kna ; Prelovsek:2019ywc ; Bulava:2019iut ; Bicudo:2021qxj ; Sadl:2021bme . The results show that the static energies are mostly flat lines at the energy corresponding to the heavy meson masses except in some cases in the short-distance limit where the potential is attractive or repulsive depending on the specific heavy-meson pair. However, only in Refs. Bali:2005fu ; Bulava:2019iut the coupled system of the heavy meson-antimeson pair and quarkonium static energies has been studied. The analysis of Ref. TarrusCastella:2022rxb indicates that when the coupling is taken into account the static potential of the heavy meson-antimeson pair becomes flat. Hence, we are going to assume that
@@
with
the wave function solution of the following coupled Schrödinger equations
@@
and $n$ is the principal quantum number. The components of the tensor spherical harmonics Khersonskii:1988krb defined as
@@
Let us compute the continuum states contribution to the double-heavy hadron self-energy
@@
where $p(E)=\sqrt{2\mu(E+i\epsilon+2m_{Q}-m_{T})}$ with $\mu$ and $m_{T}$ being the heavy hadron pair reduced and total masses, respectively. In order to take into account the self-energy to all orders in perturbation theory, we can resum the self-energy into the double-heavy hadron propagator. Using Sokhotsky’s formula we separate the real and imaginary parts of the self-energy and find the bound state energy of the double-heavy hadron as the roots of
@@
For the rest of the paper we will use $E_{n\ell}=E^{(0)}_{n\ell}$. However, one of the strengths of our approach is that we can improve the accuracy of this quantity by including higher order terms in BOEFT. This is important, because when $E_{n\ell}+2m_{Q}-m_{T}$ is close to zero, i.e when the double-heavy hadron state is close to a heavy hadron pair threshold, the contribution of the continuum states is very sensitive to small variations of $E_{n\ell}$.
The conventional quarkonium case, which was studied in Ref. TarrusCastella:2022rxb , corresponds to setting $\kappa^{pc}=0^{++}$ in the expressions of this section. In this case $\ell$ must be equal to $l$. Furthermore, $\lambda=0$ as it is the projection of $\kappa$, which corresponds to the $\Sigma_{g}^{+}$ representation of the static energy of conventional quarkonium. The wave function $\phi^{(\kappa\lambda)}_{n\ell}(r)$ in Eq. (18) reduces to the radial wave function of conventional quarkonium. To match Eq. (19)-(20) with $\kappa^{pc}=0^{++}$ to the results of Ref. TarrusCastella:2022rxb a $(-1)^{\kappa^{\prime}}$ phase has to be shifted from the definition of Eq. (19) to Eq. (20).
3 Hybrid quarkonium and heavy-meson pairs
Let us now apply the results of the previous section to study the effect of the heavy meson-antimeson thresholds to the hybrid quarkonium spectrum. In table 1 we show the quantum numbers of the four lowest mass gluelumps Herr:2023xwg and the $D_{\infty h}$ representations of the corresponding static states Brambilla:1999xf ; Berwein:2015vca . Each heavy meson-antimeson pair is characterized by the spin and parity of the light-quark states forming the two heavy mesons, which we label as $\kappa_{1}^{p_{1}}$ and $\kappa_{2}^{p_{2}}$ in table 2. Combining the spin and parity of the light quarks we find their total $\kappa^{pc}$. From these, we work out the possible $\Lambda^{\sigma}_{\eta}$ representations in table 2. A hybrid (or conventional) quarkonium state couples to a heavy meson-antimeson pair at leading order in the heavy-quark mass expansion if they can be projected into the same $D_{\infty h}$ representation. For the lowest lying hybrid quarkonium, with $\kappa=1^{+-}$ gluelump, the only matching $D_{\infty h}$ representation with the lowest lying heavy meson-antimeson pairs, i.e. $D_{(s)}^{(*)}\bar{D}_{(s)}^{(*)}(B_{(s)}^{(*)}\bar{B}_{(s)}^{(*)})$, is $\Sigma_{u}^{-}$ corresponding to a $0^{-+}$ light-quark state. For this case the form factor in Eq. (20) takes the following simple form
@@ 3 Hybrid quarkonium and heavy-meson pairs
Table 2: Total spin ($\kappa$), parity ($p$) and charge conjugation ($c$) of the light-quark state of heavy meson-antimeson pairs and their projections into $D_{\infty h}$ representations. The $\kappa^{pc}$ are obtained as all the possible combinations of the spin of each heavy meson light-quark state. Note that, the representation corresponding to the light-quark state of the heavy antimeson in the pair should be taken conjugated. The ground state heavy meson doublet corresponds to $(1/2)^{+}$ while $(1/2)^{-}$ and $(3/2)^{-}$ correspond to excited heavy meson states.
We can infer the following selection rules for the decays of hybrid quarkonium into $D_{(s)}^{(*)}\bar{D}_{(s)}^{(*)}(B_{(s)}^{(*)}\bar{B}_{(s)}^{(*)})$ from Eq. (25). First, we can see that only states with a $\lambda=0$ component can decay through this channel, as has been recently pointed out in Ref. Bruschini:2023tmm . In table 3 we summarize the details of the $\kappa=1^{+-}$ gluelump hybrid quarkonium states Berwein:2015vca ; Oncala:2017hop . Since the $\lambda=0$ component corresponds to the $\Sigma_{u}^{-}$ static state, we can read off table 3 which hybrid states are allowed to decay through this channel. Using the notation $l_{\ell}$, these are $s_{0}$, $(s\backslash d)_{1}$ and $(p\backslash f)_{2}$. On the other hand the states $p_{1}$ and $d_{2}$ are forbidden to decay. Higher angular momentum, not shown in table 3, with mixed $l$ are also allowed to decay. It is interesting to note that the mixed states are dominated by the $\Pi_{u}$ component Berwein:2015vca , which adds a small suppression to the decays into heavy mesons. The second selection rule is that the partial wave of the heavy-meson pair must be equal to the total angular momentum of the hybrid quarkonium states $\ell$. This is a consequence of the light-quark state of the heavy-meson pairs coupling to hybrids being spin 00. Therefore, only one partial wave is allowed in the decays of quarkonium hybrids into $D^{(*)}\bar{D}^{(*)}(B^{(*)}\bar{B}^{(*)})$ at leading order in BOEFT.
In previous works Tanimoto:1982eh ; LeYaouanc:1984gh ; Iddir:1988jc ; Ishida:1991mx ; Close:1994hc ; Page:1996rj ; Page:1998gz ; Kou:2005gt , it had been argued that quarkonium hybrids cannot decay into $S$-wave heavy-meson pairs, while the more recent Ref. Bruschini:2023tmm argued that these decays are not suppressed. Our results show that, in BOEFT, decays into $S$-wave heavy-meson pairs are $1/m_{Q}$-suppressed except for $\ell=0$ quarkonium hybrids.
@@ 00 $s$ $\Sigma_{u}^{-}$$\{0^{++},1^{+-}\}$ $H_{3}$
Table 3: Quantum numbers of hybrid quarkonium with $1^{+-}$ gluelump. The quantum numbers are as follows: $l(l+1)$ is the eigenvalue of $\bm{L}^{2}_{Q\bar{Q}}$, the square of the angular momentum of the heavy-quark pair. $\ell(\ell+1)$ is the eigenvalue of $\bm{L}^{2}=(\bm{L}_{Q\bar{Q}}+\bm{S}_{1})^{2}$, with $\bm{S}_{1}$ the gluelump spin operator. $s_{Q\bar{Q}}(s_{Q\bar{Q}}+1)$ is the eigenvalue of $\bm{S}^{2}_{Q\bar{Q}}$, the square of the total heavy-quark spin operator. The square of the total spin $\bm{J}^{2}=(\bm{L}+\bm{S}_{Q\bar{Q}})^{2}$ has eigenvalue $j(j+1)$. As we are working at leading order in the heavy-quark mass expansion the hybrid quarkonium masses only depend on $\ell$ and a principal quantum number $n$. The set of total spin states for each $\ell$ form degenerate multiplets at this order. The $\Lambda^{\sigma}_{\eta}$ column indicates the $D_{\infty h}$ representations of static energies that contribute to the states. For $\ell>0$, two sets of solutions with opposite parity exists. One of these corresponds to the mixing the $\Sigma_{u}^{-}$ and $\Pi_{u}$ static states while the other is purely $\Pi_{u}$erwein:2015vca . The solutions that mix static states also mix the heavy-quark pair angular momentum Oncala:2017hop . The last column indicates the labels for the spin-symmetry multiplets used in Refs. Braaten:2014qka ; Berwein:2015vca .
@@
Figure 1: Leading order diagrams in weakly-coupled pNQRCD contributing to the hybrid quarkonium and heavy-meson pair coupling potential. The diagrams are in coordinate space. The single, double and curly lines represent the heavy-quark singlet, heavy-quark octet and gluon fields respectively. The dashed line corresponds to the light-quark fields. The black dots stand for the dipolar chromoelectric operators. The circles with a cross and with a diamond represent the quarkonium hybrid and heavy meson-antimeson pair states, respectively. It should be noted that while at leading order the quarkonium hybrid only overlaps with the octet field, the heavy-meson pair can overlap both with the singlet and octet fields. The gray circles represent nonperturbative dynamics of the light degrees of freedom.
The short-distance region is, by construction, the region where the multipole expansion is valid. Therefore, in this region the potential can be expanded in a series in powers of $r$. Such series can be obtained directly using weakly-coupled pNRCQD Pineda:1997bj ; Brambilla:1999xf , which is an EFT incorporating the heavy-quark mass and multipole expansions. Several similar examples can be found in Refs. Oncala:2017hop ; Brambilla:2019jfi ; Soto:2021cgk . This computation yields the coefficients of the expanded potential as correlators of gluon and/or light-quark operators, which encode the nonperturbative dynamics. The two weakly-coupled pNRCQD leading order diagrams contributing to coupling potential are depicted in Fig. 1. We must consider two diagrams because the heavy meason-antimeson pair state can overlap both with the heavy-quark singlet and octet fields. However, in both instances, a single dipolar chromoelectric operator is enough to produce the transition between the quarkonium hybrid and heavy-meson pair. The explicit form of this operators can be found, for instance, in Eq. (3) of Ref. Brambilla:1999xf . The singlet-to-octet coupling reads as ${\rm Tr}[{\rm S}^{\dagger}\bm{r}\cdot\bm{{\rm E}}{\rm O}+h.c]$ and the octet-to-octet as ${\rm Tr}[{\rm O}^{\dagger}\bm{r}\cdot\bm{{\rm E}}{\rm O}+h.c]$. Therefore, at leading order, the short-distance coupling potential must be proportional to $r$:
@@
The long distance quarkonium-heavy-meson pair mixing potential obtained from lattice data was found to behave as $n_{l}=3$ TarrusCastella:2022rxb .
@@
$V_{01\Sigma}(r)=w_{s}(r)V^{\,{\rm(s.d.)}}_{01\Sigma}(r)+w_{l}(r)V^{\,{\rm(l.d.)}}_{01\Sigma}(r)\,,$ (29)
with the interpolating functions $w_{s}=(r_{0}/(r+r_{0}))^{n_{l}+2}$ and $w_{l}=(r/(r+r_{0}))^{n_{l}+2}$. The $r_{0}$ parameter determines the value of $r$ where both interpolating functions are equal. We will use values of $r_{0}\sim 0.25$ fm since it was found in Ref. TarrusCastella:2022rxb to provide a good description of the quarkonium-heavy-meson pair mixing potential obtained from the lattice data of Ref. Bali:2005fu as the lattice mixing potential reaches its maximum absolute value for distances $r\sim 0.25$ fm. Moreover, it is a reasonable value for the distance in which the multipole expansion breaks down and thus for the transition from the short-distance regime to the long-distance one.
We consider three different variations of the model potential. Models (i) and (ii) take the parameter values given by dimensional analysis in Eq. (28) and $n_{l}=3$ and $n_{l}=2$, respectively. Finally, model (iii) is tuned to roughly match the lattice quarkonium-heavy-meson pair mixing potential to serve as a reference point. Its parameters read as $c_{s}=8\Lambda^{2}_{\rm QCD}$, $c_{l}=0.9\Lambda^{-2}_{\rm QCD}$, $n_{l}=3$ and $r_{0}=0.275$ fm. In Fig. 2 we plot the three model potentials and the quarkonium-heavy-meson pair mixing potential fitted in Ref. TarrusCastella:2022rxb to the lattice data of Ref. Bali:2005fu . The heavy meson-antimeson pair interpolating operator in Ref. Bali:2005fu is isospin $I=0$ corresponding to the normalized sum of the charged and neutral heavy meson-antimeson pairs. As we use this data as reference, we should also take ${\cal M}_{0^{-+}}$ to be isospin $I=0$ and therefore the results obtained from this model potential will correspond to the sum of neutral and charged heavy-meson pair contributions. Since the light-quarks in Ref. Bali:2005fu are degenerate it is also possible to take the field ${\cal M}_{0^{-+}}$ as carrying a single $q\bar{q}$ light-quark flavor and rescale the potential by $1/\sqrt{2}$. We can make use of this to obtain the coupling potential to $D_{s}^{(*)}\bar{D}_{s}^{(*)}(B_{s}^{(*)}\bar{B}_{s}^{(*)})$ in the $SU(3)$ flavor symmetry limit.
Figure 2: Plot of the three models for the quarkonium hybrid and heavy-meson pair coupling potential. The short-distance parametrization is $\sim r$ as determined from weakly-coupled pNRQCD. In the first two models the long-distance pieces behave as $r^{-3}$ and $r^{-2}$, respectively, while all the parameters values are determined by dimensional analysis. In the third model the parameters are tuned, within natural sizes, to match the quarkonium and heavy meson-antimeson pair coupling potential obtained from the lattice computation of Ref. Bali:2005fu , which we use as a reference for our models. The latter corresponds to the line plotted in light gray.